OpenStudy (anonymous):

d/dx (x^2+7)^5 ( 2 +5x^2 -3x^3) i need full solution using product rule. My answer is perhaps wrong.

OpenStudy (lgbasallote):

what's your answer?

OpenStudy (anonymous):

(x^2+7) (10+70x-42x^2-5x^3-9x^4) Checked using wolframalpha it is even more complex than the actual answer answer (x^2+7)^4 (90x-63x^2+60x^3-39x^4)

OpenStudy (anonymous):

*(x^2+7)^4

OpenStudy (cornzyblack):

Ok fine this involves Product rule and Product rule goes like y = uv then, dy/dx = u* (dv/dx) + v*(du/dx) ok so in this question we take y = (x^2 + 7)^5 (2 + 5x^2 - 3x^3) So our u = (x^2 + 7) ^ 5 and our v = (2 + 5x^2 - 3x^3) using Product rule (that is uv = (u*( dv/dx) + v*(du/dx)) so our du/dx = 5(x^2 + 7)^4 * (2x) = 10x(x^2 + 7)^4 and our dv/dx = 10x - 9x^2 Substituting into product rule we have (x^2 + 7)^5 (10x - 9x^2) + (2 + 5x^ 2 - 3x^3)(10x(x^2 +7)^4) And i stop here :)